OpenStudy (anonymous):

A bike accelerates uniformly from rest to a speed of 7.10 m/s over a distance of 25.4 m. Determine the acceleration of the bike.

OpenStudy (anonymous):

so it would be helpful to write down what it given. and then you can see which formula you can use.

OpenStudy (anonymous):

initial velocity= Vi= 0 m/s final velocity= Vf= 7.10m/s distance= S= 25.4m use formula, 2aS= Vf^2 - Vi^2 can you solve it now?

OpenStudy (ksaimouli):